in ydse,d=2mm,D=2m,lambda=500nm,if intensity of 2 slits are Io and 9Io,then find the intensity at x=1/6 mm from central maxima.

Grade:12

2 Answers

Arun
6 years ago
Given:
d=2mm
D=2m  
lambda=500nm
Intensities : Io & 9Io
path difference = xd/D
                        = 1/6 * 10 -3 * 2 * 10 -3   /   2
                       = 1.66 * 10 -7 m
phase Difference =  2 π   ∆ x   /   λ =   2 . 08   =   0 . 664   π   =   119 . 52   d e g r e e
therefore 
I net = Io + 9Io + 2 (9Io2)1/2 cos (119.52)
       = 10 Io - 2.9Io
       = 7.04 Io   ans
Yash Chourasiya
3 years ago
Dear Student

Given:
d = 2mm
D = 2m
lambda = 500nm
Intensities : Io & 9Io
path difference = xd/D
= 1/6 * 10-3* 2 * 10-3 / 2
= 1.66 * 10-7m
phase Difference =2 π ∆ x / λ = 2 . 08 = 0 . 664 π = 119 . 52 d e g r e e
therefore
I net = Io + 9Io+ 2 (9Io2)1/2cos(119.52)
= 10 Io- 2.9Io
= 7.04 Io

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya